# Intuitionistic NF set theory

### v5

@@ 11. Cartesian products
Proof. This follows in a few steps from extensionality, the definition of $\times$, and the logical fact that
@@
Let $Y$ be a finite set and let $a$ be any set. Then $\{a\}\times Y$ is finite. If $\kappa=Nc(Y)$ then ${\mathbb{T}}^{2}\kappa=Nc(\{a\}\times Y)$.
Proof. Consider the map $f:USC^{2}(Y)\to\{a\}\times Y$ defined by
@@
The formula is stratified, giving $y$ and $a$ index 0, so $\langle a,y\rangle$ gets index 2, as does $\{\{y\}\}$. $Y$ gets index 1. Since the formula is stratified, $f$ can be defined in INF.
One then proves without any surprises that $f$ is a similarity from $USC^{2}(Y)$ to $a\times Y$. We omit the straightforward 196-line verification of that fact.
@@
$USC^{2}(Y)\sim\{a\}\times Y$   since $f$ is a similarity
$USC(Y)\in{\sf FINITE}$   by Lemma 3.10
$USC^{2}(Y)\in{\sf FINITE}$   by Lemma 3.10
$Y\in{\sf FINITE}$   by Lemma 4.23
$Nc(USC^{2}(Y)={\mathbb{T}}^{2}\kappa$   by definition of ${\mathbb{T}}$
$Nc(Y)={\mathbb{T}}^{2}\kappa$   by Lemma 4.8
@@
Let $X$ and $Y$ be finite sets. Then $X\times Y$ is finite. Moreover, if $\kappa=Nc(X)$ and $\mu=Nc(Y)$, then
${\mathbb{T}}^{2}(\kappa)\cdot{\mathbb{T}}^{2}(\mu)=Nc(X\times Y).$
@@
Proof. The formula to be proved is
@@
That formula (and the hypotheses listed before it) are stratified, giving $X$ and $Y$ index 1; then $X\times Y$ gets index 3, $Nc(X\times Y)$ gets index 4, $\kappa=Nc(X)$ gets index 2, and ${\mathbb{T}}^{2}(\kappa)$ gets index 4; since multiplication is a function, the whole left-hand side gets index $4$. ${\sf FINITE}$ is just parameter. Therefore we may proceed by induction on finite sets $X$.
Base case. We have to show $\Lambda\times Y\in{\sf FINITE}$. One shows $\Lambda\times Y=\Lambda$ using the definition of $\times$, and then $\Lambda\in{\sf FINITE}$ by Lemma 3.6.
@@
(73) $\forall Y\in{\sf FINITE}\,(X\times Y\in{\sf FINITE})$
@@
$(X\times Y)\cap(\{a\}\times Y)=\Lambda$   since $a\not\in X$
@@
That completes the induction step. That completes the proof of the lemma.
@@
Let $X$ and $Y$ be finite sets. If $\kappa=Nc(X)$ and $\mu=Nc(Y)$, then
${\mathbb{T}}^{2}(\kappa)\cdot{\mathbb{T}}^{2}(\mu)=Nc(X\times Y).$
@@
Proof. By induction on finite sets, like Lemma 11.3. We omit the proof, since we never use this lemma. It is included only because it illustrates the general situation that arises from using Kuratowski pairing, which increases the type.
12. Onto and one-to-one for maps between finite sets
@@
$Rel(f)$ means that all the members of $f$ are ordered pairs. $f\in{\sf FUNC}$ means that two ordered pairs in $f$ with the same first member have the same second member. (Nothing is said about possible members of $f$ that are not ordered pairs.) $f:X\to Y$ means that if $x\in X$, there is a unique $y$ such that $\langle x,y\rangle\in f$ and that $y$ is in $Y$. (But nothing is said about $\langle x,y\rangle\in f$ with $x\not\in X$.) “$f$ is one-to-one from $X$ to $Y$”, or $oneone(f,X,Y)$, means $f:X\to Y$ and in addition, if $\langle x,y\rangle\in f$ and $\langle u,y\rangle\in f$ then $x=u$, and if $y\in Y$ then $x\in X$. (So $x=u$ does not require $y\in Y$ or $x\in X$.) In particular, $f:X\to Y$ does not require $domX\subseteq X$, so the identity function maps $X$ to $X$ for every $X$; but the identity function (on the universe) has to be restricted to $X$ before it is one-to-one.
@@
Proof. By induction on finite sets $A$ we prove that for all finite sets $B$, if the domain of $f$ is $A$ and $f:A\to B$, then $f$ is finite.
Base case. A function with domain $\Lambda$ is the empty function, which is finite.
@@
Let $g:=f-\{\langle c,y\rangle\}$. One can verify that $g:A\to B$ and the domain of $g$ is $A$ [^6] Formalizing this sort of lemma makes one appreciate the informal functional notation; this lemma took 330 lines of Lean and several hours. I changed “One can easily verify” to the present “One can verify.” Then by the induction hypothesis, $g$ is finite. Since $A$ and $B$ are finite, equality on $A$ and $B$ is decidable, so any member of $f$ is either equal to $\langle c,y\rangle$ or not. Therefore
@@
Since $g$ is finite and $\{\langle c,y\rangle\}\not\in f$, $f$ is also finite. That completes the proof of the lemma.
@@
$Z=\Lambda\ \lor\ \exists x\,(x\in Z)$   by Lemma 3.4
Putting in the definition of $Z$, we have the formula in the conclusion of the lemma. That completes the proof of the lemma.
@@
Proof. By induction on finite sets, we prove that if $f:X\to X$ is one-to-one, then $f$ is onto. By Lemma 3.3, $X$ has decidable equality.
@@
(74) $\forall y\in X\,\exists x\in X\,(\langle x,y\rangle\in f)$
@@
(75) $\exists y\,(\langle t,y\rangle\in g){\ \leftrightarrow\ }t\in B$
@@
Case 2, $\neg\,\exists x\in X\,(\langle x,a\rangle\in f)$. Let $g$ be $f$ restricted to $B$. Then $Rel(g)$, and $dom(g)=B$, and $g$ is one-to-one, and $g:B\to B$. Then by the induction hypothesis, $g$ is onto. Since $f:X\to X$, there exists some $b\in X$ such that $\langle a,b\rangle\in f$. By hypothesis $b\neq a$. Then $b\in B$. Since $g$ is onto, there exists $x\in B$ such that $\langle x,b\rangle\in g$. Then $\langle x,b\rangle\in f$. Since $f$ is one-to-one, we have $x=a$. But $x\in B$, while $a\not\in B$. That contradiction completes Case 2. That completes the proof of the theorem.
@@
Let $B\in{\sf FINITE}$ and $a\not\in B$. Then $Nc(B\cup\{a\})=(Nc(B))^{+}$.
Proof. We have
$B\in Nc(B)$   by Lemma 4.11
$B\cup\{a\}\in Nc(B\cup\{a\})$   by Lemma 4.11
$B\cup\{a\}\in(NcB)^{+}$   by definition of successor
@@
$Nc(B\cup\{a\}\in{\mathbb{F}}$   by Lemma 4.21
$Nc(B)\in{\mathbb{F}}$   by Lemma 4.21
$(Nc(B))^{+}\in{\mathbb{F}}$   by Lemma 4.19
$B\cup\{a\}\in Nc(B\cup\{a\})\cap(Nc(B))^{+}$   by the definition of intersection
$Nc(B\cup\{a\}=(Nc(B))^{+}$   by Lemma 5.7
That completes the proof of the lemma.
@@
$Nc(Z)\leq Nc(X).$
Proof. We have
$Nc(X)\in{\mathbb{F}}$   by Lemma 4.21
@@
$Nc(Z)\in{\mathbb{F}}$   by Lemma 4.21
$X\in Nc(X)$   by Lemma 4.11
$Z\in Nc(Z)$   by Lemma 4.11
$Nc(Z)\leq Nc(X)$   by the definition of $\leq$
That completes the proof of the lemma.
@@
Proof. We prove the more general fact that if $X$ and $Y$ are finite sets with $Nc(X)\leq Nc(Y)$, and $f:X\to Y$ is onto, then $f$ is one-to-one. (The theorem follows by taking $Y=X$). More explicitly, we will prove by induction on finite sets $Y$ that
$\forall Y\in{\sf FINITE}\,\forall X\in{\sf FINITE}\,(Nc(X)\leq Nc(Y)\ {\rightarrow}\ \forall f\,(f\in{\sf FUNC}$
@@
The formula is stratified, giving $x,y,z$ index 0, $f$ index 3, $X$ and $Y$ index 1, and $Nc(X)$ and $Nc(Y)$ index 2. ${\sf FUNC}$ and ${\sf FINITE}$ are parameters; $Rel(f)$ is stratified giving $f$ index 3; $dom(f)\subseteq X$ can be expressed as $\forall x,y\,(\langle x,y\rangle\in f\ {\rightarrow}\ x\in X)$, which is stratified. Therefore we may proceed by induction on finite sets $Y$.
Base case, $Y=\Lambda$. Then (in the last line) $y\in Y$ is impossible, so the last line holds if the previous lines are assumed. That completes the base case.
@@
(77) $\forall x\in X\,(\langle x,a\rangle\in f\ \lor\ \langle x,a\rangle\not\in f).$
To prove that, let $x\in X$. Since $f:X\to Y$, there exists $y\in Y$ with $\langle x,y\rangle\in f$. Since $f\in{\sf FUNC}$, we have $\langle x,a\rangle\in f{\ \leftrightarrow\ }y=a$. Since $Y$ is finite, we have $y=a\ \lor y\neq a$ by Lemma 3.3. That completes the proof of (77). Then by Lemma 3.19, $Z\in{\sf FINITE}$ and $X-Z\in{\sf FINITE}$.
Let $g$ be $f$ restricted to $X-Z$. Then $g:X-Z\to B$ and $g$ is onto $B$. I say that
(78) $Nc(X-Z)\neq Nc(X)$
To prove that, assume $Nc(X-Z)=Nc(X)$. Then
$Nc(X-Z)\in{\mathbb{F}}$   by Lemma 4.21
$Nc(X)\in{\mathbb{F}}$   by Lemma 4.21
@@
Now I say that $Nc(X-Z)\leq Nc(B)$. To prove that:
$Nc(X-Z)\leq Nc(X)$   by Lemma 12.7
$Nc(X-Z)<Nc(X)$   by (78 and the definition of $<$
$Nc(X)\leq Nc(B\cup\{a\})$   by hypothesis
$Nc(B\cup\{a\})=(Nc(B))^{+}$   since $a\not\in B$
$Nc(X-Z)<Nc(B)^{+}$   by the previous two lines
$Nc(X-Z)\leq Nc(B)$   by Lemma 12.6
@@
$Nc(X-Z)=Nc(B)$ by Lemma 4.9 and ten omitted steps
$Nc(X)=Nc(X-Z)+Nc(Z)$ by Lemma 8.15
$Nc(X)=Nc(B)+Nc(Z)$ by the previous two lines
$Nc(X)\leq Nc(Y)$ by hypothesis
$Nc(B)+Nc(Z)\leq Nc(Y)$ by the previous two lines
$Nc(Y)=Nc(B)^{+}$ since $Y=B\cup\{a\}$ and $a\not\in B$
$Nc(B)+Nc(Z)\leq Nc(B)^{+}$ by the previous two lines
$Nc(Z)={\sf one}$ by Lemma 12.6
@@
Case 2, $y\neq a$. Then $u\not\in Z$ and $v\not\in Z$, so $\langle u,y\rangle\in g$ and $\langle v,y\rangle\in g$. Since $g$ is one-to-one, we have $u=v$ as desired. That completes Case 2. That completes the induction step. That completes the proof of the lemma.
@@
Let $X$ and $Y$ be finite sets, and suppose $f:X\to Y$ is onto, and the domain of $f$ is $X$. Then $Nc(Y)\leq Nc(X)$.
Proof. By induction on finite sets $X$, we prove the theorem for all $Y$.
Base case. If $f:\Lambda\to Y$ has domain $\Lambda$ and is onto $Y$ then $Y=\Lambda$, so
$Nc(X)=Nc(Y)=Nc(\Lambda)={\sf zero}.$
@@
$Nc(Y)\leq Nc(X)$   by the induction hypothesis
$Nc(X)<Nc(X)^{+}$   by Lemma 5.25
$Nc(X)^{+}=Nc(X\cup\{c\})$   by Lemma 4.13
$Nc(Y)\leq Nc(X\cup\{c\})$   by the preceding lines
@@
$\exists u\,(u\in Nc(X)^{+})$   namely $u=X\cup\{c\}$
$\exists u\,(u\in Nc(Y-\{t\})^{+})$   namely $u=(Y-\{t\})\cup\{t\}=Y$
$Nc(X)^{+}\in{\mathbb{F}}$   by Lemma 4.19
$Nc(Y-\{t\})^{+}\in{\mathbb{F}}$   by Lemma 4.19
@@
(79) $\hskip-36.98866ptNc(Y-\{t\})^{+}\leq Nc(X)^{+}$   by Lemma 5.11
$Y\in Nc(Y)$   by Lemma 4.11
$(Y-\{t\})\cup\{t\}\in Nc(Y-\{t\})^{+}$   by definition of successor
$Y\in Nc(Y-\{t\})^{+}$   since $(Y-\{t\})\cup\{t\}=Y$
$Nc(Y-\{t\})^{+}=Nc(Y)$   by Lemma 5.7
$Nc(Y)\leq Nc(X)^{+}$   by (79) and the preceding line
$Nc(X\cup\{c\})=Nc(X)^{+}$   by Lemma 4.13
$Nc(Y)\leq Nc(X\cup\{c\})$   since $Nc(Y)\leq Nc(X)^{+}=Nc(X\cup\{c\})$
That completes the induction step. That completes the proof of the lemma.
@@
Let $A$ and $B$ be finite sets, and let $f$ be a function mapping $A$ onto $B$. Then $Nc(B)\leq Nc(A).$
Proof. We may assume without loss of generality that $A$ is the domain of $f$. Then
@@
(80) $\forall t\in B\,(t\in Z\ \lor\ t\not\in Z)$
Now we will proceed by induction on finite sets $A$ to prove that for all finite sets $B$ and all $g:A\to B$ onto, $Nc(B)\leq Nc(A)$.
Base case: If $g:\Lambda\to B$ is onto, then $B=\Lambda$, so $Nc(A)=Nc(B)=Nc(\Lambda)$.
Induction step: Let $g:A\cup\{c\}\to B$ be onto, where $c\not\in A$. Let $t=g(c)$ and let $f=g-\{\langle c,t\rangle\}$. Then $f:A\to B$. By(80, $t\in Z\ \lor\ t\not\in Z$. That is,
@@
Case 2. If there does not exist such a $b$ then $f:A\to(B-\{t\})$ is onto. Also $B-\{t\}$ is a finite set, by Lemma 3.20. Hence by the induction hypothesis, $Nc(B-\{t\})\leq Nc(A)$. Then
@@
$Nc(B)=(Nc(B-\{t\})^{+}$
$(Nc(B-\{t\})^{+}\leq(Nc(A))^{+}$   by Lemma 5.11
$Nc(B)\leq(Nc(A))^{+}$   by the previous two lines
$Nc(A)^{+}=Nc(A\cup\{c\})$   by Lemma 4.13, since $c\not\in A$
$Nc(B)\leq Nc(A\cup\{c\})$   by the previous two lines
That completes Case 2, and that completes the induction step, and that completes the proof of the lemma.
@@
Proof. We have
$a\in SSC(X)$   by Lemma 3.18
$b\in SSC(X)$   by Lemma 3.18
$\forall x\in X\,(x\in a\ \lor\ x\not\in a)$   by the definition of $SSC(X)$
$\forall x\in X\,(x\in b\ \lor\ x\not\in b)$   by the definition of $SSC(X)$
@@
Let $X$ be a finite set and let $y$ be a finite subset of $SSC(X)$ (that is, the members of $y$ are separable subsets of $X$). Then the union of $y$ is a finite set. That is,
@@
Proof. By induction on finite sets $y$ (for fixed $X$).
Base case. When $y=\Lambda$, the union of $y$ is also $\Lambda$, which is finite.
Induction step. Suppose $c\not\in y$ and $y\cup\{c\}\subseteq SSC(X)$. Then we have (in a few steps from the definitions of $\bigcup$ and $\cup$)
(81) $\bigcup\,(y\cup\{c\})=\left(\bigcup y\right)\cup c$
@@
$c\in SSC(X)$   since $y\cup\{c\}\subseteq SSC(X)$
@@
$y\subseteq SSC(X)$   since $y\cup\{c\}\subseteq SSC(X)$
$\bigcup y\subseteq X$   since $y\subseteq SSC(X)$

### v6

@@ 11 Cartesian products
Proof.
This follows in a few steps from extensionality, the definition of $\times$, and the logical fact that
@@
Let $Y$ be a finite set and let $a$ be any set. Then $\{a\}\times Y$ is finite. If $\kappa=|Y|$ then ${\mathbb{T}}^{2}\kappa=|\{a\}\times Y|$.
Proof.
Consider the map $f:\mathcal{P}_{1}^{2}(Y)\to\{a\}\times Y$ defined by
@@
The formula is stratified, giving $y$ and $a$ index 0, so $\langle a,y\rangle$ gets index 2, as does $\{\{y\}\}$. $Y$ gets index 1. Since the formula is stratified, $f$ can be defined in $\mathit{i}\mkern 0.4mu\mathrm{NF}$.
One then proves without any surprises that $f$ is a similarity from $\mathcal{P}_{1}^{2}(Y)$ to $\{a\}\times Y$. We omit the straightforward 196-line verification of that fact.
@@
$\mathcal{P}_{1}^{2}(Y)\sim\{a\}\times Y$   since $f$ is a similarity
$\mathcal{P}_{1}(Y)\in{\sf FINITE}$   by Lemma 3.10
$\mathcal{P}_{1}^{2}(Y)\in{\sf FINITE}$   by Lemma 3.10
$\{a\}\times Y\in{\sf FINITE}$   by Lemma 4.23
$|\mathcal{P}_{1}^{2}(Y)|={\mathbb{T}}^{2}\kappa$   by definition of $\mathbb{T}$}
$|\{a\}\times Y|={\mathbb{T}}^{2}\kappa$   by Lemma 4.8
@@
Let $X$ and $Y$ be finite sets. Then $X\times Y$ is finite. Moreover, if $\kappa=|X|$ and $\mu=|Y|$, then
${\mathbb{T}}^{2}(\kappa)\cdot{\mathbb{T}}^{2}(\mu)=|X\times Y|.$
@@
Proof.
The formula to be proved is
@@
That formula (and the hypotheses listed before it) are stratified, giving $X$ and $Y$ index 1; then $X\times Y$ gets index 3, $|X\times Y|$ gets index 4, $\kappa=|X|$ gets index 2, and ${\mathbb{T}}^{2}(\kappa)$ gets index 4; since multiplication is a function, the whole left-hand side gets index $4$. $\sf FINITE$} is just parameter. Therefore we may proceed by induction on finite sets $X$.
Base case. We have to show $\emptyset\times Y\in{\sf FINITE}$. One shows $\emptyset\times Y=\emptyset$ using the definition of $\times$, and then $\emptyset\in{\sf FINITE}$ by Lemma 3.6.
@@
$\forall Y\in{\sf FINITE}\,(X\times Y\in{\sf FINITE})$ (80)
@@
$(X\times Y)\cap(\{a\}\times Y)=\emptyset$   since $a\not\in X$
@@
That completes the induction step. ∎
@@
Let $X$ and $Y$ be finite sets. If $\kappa=|X|$ and $\mu=|Y|$, then
${\mathbb{T}}^{2}(\kappa)\cdot{\mathbb{T}}^{2}(\mu)=|X\times Y|.$
@@
Proof.
By induction on finite sets, like Lemma 11.3. We omit the proof, since we never use this lemma. It is included only because it illustrates the general situation that arises from using Kuratowski pairing, which increases the type. ∎
12 Onto and one-to-one for maps between finite sets
@@
$Rel(f)$ means that all the members of $f$ are ordered pairs. $f\in{\sf FUNC}$ means that two ordered pairs in $f$ with the same first member have the same second member. (Nothing is said about possible members of $f$ that are not ordered pairs.) $f:X\to Y$ means that if $x\in X$, there is a unique $y$ such that $\langle x,y\rangle\in f$ and that $y$ is in $Y$. (But nothing is said about $\langle x,y\rangle\in f$ with $x\not\in X$.) “$f$ is one-to-one from $X$ to YY”, or $oneone(f,X,Y)$, means $f:X\to Y$ and in addition, if $\langle x,y\rangle\in f$ and $\langle u,y\rangle\in f$ then $x=u$, and if $y\in Y$ then $x\in X$. (So $x=u$ does not require $y\in Y$ or $x\in X$.) In particular, $f:X\to Y$ does not require $domX\subseteq X$, so the identity function maps $X$ to $X$ for every $X$; but the identity function (on the universe) has to be restricted to $X$ before it is one-to-one.
@@
Proof.
By induction on finite sets $A$ we prove that for all finite sets $B$, if the domain of $f$ is $A$ and $f:A\to B$, then $f$ is finite.
Base case. A function with domain $\emptyset$ is the empty function, which is finite.
@@
Let $g:=f-\{\langle c,y\rangle\}$. One can verify that $g:A\to B$ and the domain of $g$ is AA. [^11] Formalizing this sort of lemma makes one appreciate the informal functional notation; this lemma took 330 lines of Lean and several hours. I changed “One can easily verify” to the present “One can verify.” Then by the induction hypothesis, $g$ is finite. Since $A$ and $B$ are finite, equality on $A$ and $B$ is decidable, so any member of $f$ is either equal to $\langle c,y\rangle$ or not. Therefore
@@
Since $g$ is finite and $\{\langle c,y\rangle\}\not\in f$, $f$ is also finite. ∎
@@
$Z=\emptyset\ \lor\ \exists x\,(x\in Z)$   by Lemma 3.4
Putting in the definition of $Z$, we have the formula in the conclusion of the lemma. ∎
@@
Proof.
By induction on finite sets, we prove that if $f:X\to X$ is one-to-one, then $f$ is onto. By Lemma 3.3, $X$ has decidable equality.
@@
$\forall y\in X\,\exists x\in X\,(\langle x,y\rangle\in f)$ (81)
@@
$\exists y\,(\langle t,y\rangle\in g){\ \leftrightarrow\ }t\in B$ (82)
@@
Case 2, $\neg\,\exists x\in X\,(\langle x,a\rangle\in f)$. Let $g$ be $f$ restricted to $B$. Then $Rel(g)$, and $dom(g)=B$, and $g$ is one-to-one, and $g:B\to B$. Then by the induction hypothesis, $g$ is onto. Since $f:X\to X$, there exists some $b\in X$ such that $\langle a,b\rangle\in f$. By hypothesis $b\neq a$. Then $b\in B$. Since $g$ is onto, there exists $x\in B$ such that $\langle x,b\rangle\in g$. Then $\langle x,b\rangle\in f$. Since $f$ is one-to-one, we have $x=a$. But $x\in B$, while $a\not\in B$. That contradiction completes Case 2. ∎
@@
Let $B\in{\sf FINITE}$ and $a\not\in B$. Then $|B\cup\{a\}|=(|B|)^{+}$.
Proof.
We have
$B\in|B|$   by Lemma 4.11
$B\cup\{a\}\in|B\cup\{a\}|$   by Lemma 4.11
$B\cup\{a\}\in(|B|)^{+}$   by definition of successor
@@
$|B\cup\{a\}|\in{\mathbb{F}}$   by Lemma 4.21
$|B|\in{\mathbb{F}}$   by Lemma 4.21
$(|B|)^{+}\in{\mathbb{F}}$   by Lemma 4.19
$B\cup\{a\}\in|B\cup\{a\}|\cap(|B|)^{+}$   by the definition of intersection
$|B\cup\{a\}|=(|B|)^{+}$   by Lemma 4.24
@@
$|Z|\leq|X|.$
Proof.
We have
$|X|\in{\mathbb{F}}$   by Lemma 4.21
@@
$|Z|\in{\mathbb{F}}$   by Lemma 4.21
$X\in|X|$   by Lemma 4.11
$Z\in|Z|$   by Lemma 4.11
$|Z|\leq|X|$   by the definition of $\leq$
@@
Proof.
We prove the more general fact that if $X$ and $Y$ are finite sets with $|X|\leq|Y|$, and $f:X\to Y$ is onto, then $f$ is one-to-one. (The theorem follows by taking $Y=X$). More explicitly, we will prove by induction on finite sets $Y$ that
$\forall Y\in{\sf FINITE}\,\forall X\in{\sf FINITE}\,(|X|\leq|Y|\ {\rightarrow}\ \forall f\,(f\in{\sf FUNC}$
@@
The formula is stratified, giving $x,y,z$ index 0, $f$ index 3, $X$ and $Y$ index 1, and $|X|$ and $|Y|$ index 2. $\sf FUNC$} and $\sf FINITE$} are parameters; $Rel(f)$ is stratified giving $f$ index 3; $dom(f)\subseteq X$ can be expressed as $\forall x,y\,(\langle x,y\rangle\in f\ {\rightarrow}\ x\in X)$, which is stratified. Therefore we may proceed by induction on finite sets $Y$.
Base case, $Y=\emptyset$. Then (in the last line) $y\in Y$ is impossible, so the last line holds if the previous lines are assumed. That completes the base case.
@@
$\forall x\in X\,(\langle x,a\rangle\in f\ \lor\ \langle x,a\rangle\not\in f).$ (84)
[…]
Let $g$ be $f$ restricted to XZX-Z. Then $g:X-Z\to B$ and $g$ is onto $B$. I say that
$|X-Z|\neq|X|$ (85)
To prove that, assume |XZ|=|X||X-Z|=|X|. Then
$|X-Z|\in{\mathbb{F}}$   by Lemma 4.21
$|X|\in{\mathbb{F}}$   by Lemma 4.21
@@
Now I say that $|X-Z|\leq|B|$. To prove that:
$|X-Z|\leq|X|$   by Lemma 12.7
$|X-Z|<|X|$   by (85) and the definition of <<
$|X|\leq|B\cup\{a\}|$   by hypothesis
$|B\cup\{a\}|=(|B|)^{+}$   since $a\not\in B$
$|X-Z|<|B|^{+}$   by the previous two lines
$|X-Z|\leq|B|$   by Lemma 12.6
@@
$|X-Z|=|B|$ by Lemma 4.9 and ten omitted steps
$|X|=|X-Z|+|Z|$ by Lemma 8.15
$|X|=|B|+|Z|$ by the previous two lines
$|X|\leq|Y|$ by hypothesis
$|B|+|Z|\leq|Y|$ by the previous two lines
$|Y|=|B|^{+}$ since $Y=B\cup\{a\}$ and $a\not\in B$
$|B|+|Z|\leq|B|^{+}$ by the previous two lines
$|Z|={\sf one}$ by Lemma 12.6
@@
Case 2, $y\neq a$. Then $u\not\in Z$ and $v\not\in Z$, so $\langle u,y\rangle\in g$ and $\langle v,y\rangle\in g$. Since $g$ is one-to-one, we have $u=v$ as desired. That completes Case 2. That completes the induction step. ∎
@@
Let $X$ and $Y$ be finite sets, and suppose $f:X\to Y$ is onto, and the domain of $f$ is $X$. Then $|Y|\leq|X|$.
Proof.
By induction on finite sets $X$, we prove the theorem for all $Y$.
Base case. If $f:\emptyset\to Y$ has domain $\emptyset$ and is onto $Y$ then $Y=\emptyset$, so
$|X|=|Y|=|\emptyset|={\sf zero}.$
@@
$|Y|\leq|X|$   by the induction hypothesis
$|X|<|X|^{+}$   by Lemma 5.26
$|X|^{+}=|X\cup\{c\}|$   by Lemma 4.13
$|Y|\leq|X\cup\{c\}|$   by the preceding lines
@@
$\exists u\,(u\in|X|^{+})$   namely $u=X\cup\{c\}$
$\exists u\,(u\in|Y-\{t\}|^{+})$   namely $u=(Y-\{t\})\cup\{t\}=Y$
$|X|^{+}\in{\mathbb{F}}$   by Lemma 4.19
$|Y-\{t\}|^{+}\in{\mathbb{F}}$   by Lemma 4.19
@@
$\thinspace|Y-\{t\}|^{+}\leq|X|^{+}$   by Lemma 5.10 (86)
$Y\in|Y|$   by Lemma 4.11
$(Y-\{t\})\cup\{t\}\in|Y-\{t\}|^{+}$   by definition of successor
$Y\in|Y-\{t\}|^{+}$   since $(Y-\{t\})\cup\{t\}=Y$
$|Y-\{t\}|^{+}=|Y|$   by Lemma 4.24
$|Y|\leq|X|^{+}$   by (86) and the preceding line
$|X\cup\{c\}|=|X|^{+}$   by Lemma 4.13
$|Y|\leq|X\cup\{c\}|$   since $|Y|\leq|X|^{+}=|X\cup\{c\}|$
That completes the induction step. ∎
@@
Let $A$ and $B$ be finite sets, and let $f$ be a function mapping $A$ onto $B$. Then $|B|\leq|A|.$
Proof.
We may assume without loss of generality that $A$ is the domain of $f$. Then
@@
$\forall t\in B\,(t\in Z\ \lor\ t\not\in Z)$ (87)
Now we will proceed by induction on finite sets $A$ to prove that for all finite sets $B$ and all $g:A\to B$ onto, $|B|\leq|A|$.
Base case: If $g:\emptyset\to B$ is onto, then $B=\emptyset$, so $|A|=|B|=|\emptyset|$.
Induction step: Let $g:A\cup\{c\}\to B$ be onto, where $c\not\in A$. Let $t=g(c)$ and let $f=g-\{\langle c,t\rangle\}$. Then $f:A\to B$. By (87), $t\in Z\ \lor\ t\not\in Z$. That is,
@@
Case 2. If there does not exist such a $b$ then $f:A\to(B-\{t\})$ is onto. Also $B-\{t\}$ is a finite set, by Lemma 3.20. Hence by the induction hypothesis, $|B-\{t\}|\leq|A|$. Then
@@
$|B|=|B-\{t\}|^{+}$
$|B-\{t\}|^{+}\leq(|A|)^{+}$   by Lemma 5.10
$|B|\leq(|A|)^{+}$   by the previous two lines
$|A|^{+}=|A\cup\{c\}|$   by Lemma 4.13, since $c\not\in A$
$|B|\leq|A\cup\{c\}|$   by the previous two lines
That completes Case 2, and that completes the induction step. ∎
@@
Remark. We cannot prove the union of two finite sets is finite without some additional hypothesis, for consider $\{p\}\cup\{q\}$, where we do not know whether $p=q$ or not, e.g., $p=\emptyset$ and $q=\{x:x=\emptyset\ \land\ P\}$, where $P$ is Goldbach’s conjecture or the Riemann hypothesis. Does the union contain one or two elements?
Proof.
We have
$a\in\mathcal{P}_{s}(X)$   by Lemma 3.18
$b\in\mathcal{P}_{s}(X)$   by Lemma 3.18
$\forall x\in X\,(x\in a\ \lor\ x\not\in a)$   by the definition of $\mathcal{P}_{s}(X)$
$\forall x\in X\,(x\in b\ \lor\ x\not\in b)$   by the definition of $\mathcal{P}_{s}(X)$
@@
Let $X$ be a finite set and let $y$ be a finite subset of $\mathcal{P}_{s}(X)$ (that is, the members of $y$ are separable subsets of $X$). Then the union of $y$ is a finite set. That is,
@@
Proof.
By induction on finite sets $y$ (for fixed $X$).
Base case. When $y=\emptyset$, the union of $y$ is also $\emptyset$, which is finite.
Induction step. Suppose $c\not\in y$ and $y\cup\{c\}\subseteq\mathcal{P}_{s}(X)$. Then we have (in a few steps from the definitions of $\bigcup$ and $\cup$)
$\bigcup\,(y\cup\{c\})=\left(\bigcup y\right)\cup c$ (88)
@@
$c\in\mathcal{P}_{s}(X)$   since $y\cup\{c\}\subseteq\mathcal{P}_{s}(X)$
@@
$y\subseteq\mathcal{P}_{s}(X)$   since $y\cup\{c\}\subseteq\mathcal{P}_{s}(X)$
$\bigcup y\subseteq X$   since $y\subseteq\mathcal{P}_{s}(X)$